Please help me with this algebra problem.
Given the equation y=2/7x + 3, find the y coordinates of (-7) (0) and (7) please show work and if you can please plot the points and graph the linear equation. I just can not get this one!

Answer by rfer(16322)

y=2/7*-7+3
y=-2+3
y=1
(-2,1)
-------------
y=2/7*0+3
y=3
(0,3)
------------
y=2/7*7+3
y=2+3
y=5
(7,5)